OpenStudy (anonymous):

HELPPP FAST :( sqrt 5(8+3 sqrt 6) 8 sqrt5 + 3 sqrt30 5 sqrt 10 130 None of the above

OpenStudy (anonymous):

\[\sqrt{5}(8+3\sqrt{6)}\]

OpenStudy (anonymous):

is it the problem @tinaballerina ?

OpenStudy (anonymous):

yessum @sritama

OpenStudy (anonymous):

just multiply \[(8+3\sqrt{6}) with \sqrt{5}\]

OpenStudy (anonymous):

you will find option A is correct

OpenStudy (anonymous):

Thank you very much @sritama

OpenStudy (anonymous):

could you help with my other problem?

OpenStudy (anonymous):

welcome :) yeah,i will try..

OpenStudy (anonymous):

you seem to comprehend my questions , while everyone else is smart about their answers

OpenStudy (anonymous):

oh,is it? :)

OpenStudy (anonymous):

\[(4-\sqrt{3} )(12+5\sqrt{3)}\]

OpenStudy (anonymous):

\[(4\times12)+20\sqrt{3}-12\sqrt{3}-15\]

OpenStudy (anonymous):

the answer is 33+8 sqart 3

OpenStudy (anonymous):

ugh i need to understand this by early january for finals but , how did you learn all of this so well? do you have any ways to understand better or any websites?

OpenStudy (anonymous):

you can take the help of http://www.wolframalpha.com/

OpenStudy (anonymous):

actually all you have to do is to multiply

OpenStudy (anonymous):

if you have a problem like that \[(a+\sqrt{b})(c+\sqrt{d})\] the answer is \[ac+c \sqrt{b}+a \sqrt{d}+\sqrt{bd}\]

OpenStudy (anonymous):

OH I REMEMBER NOW.

OpenStudy (anonymous):

F. O. I. L

OpenStudy (anonymous):

Thank you so very much

OpenStudy (anonymous):

yeah..right :)

OpenStudy (anonymous):

i wanna see if i get this problem right, will you check it??

OpenStudy (anonymous):

of course..

OpenStudy (anonymous):

post it

OpenStudy (anonymous):

\[(5+\sqrt{3}) (5-\sqrt{3})\]

OpenStudy (anonymous):

\[25-12\sqrt{3}\]

OpenStudy (anonymous):

can you tell me how do you got it?

OpenStudy (anonymous):

5x5 first, equaling 25

OpenStudy (anonymous):

right,then?

OpenStudy (anonymous):

sqrt 3 times 5

OpenStudy (anonymous):

yes..next?

OpenStudy (anonymous):

sqrt 3 times - sqrt 3

OpenStudy (anonymous):

oh im sorry i skipped one

OpenStudy (anonymous):

5 times -sqrt 3

OpenStudy (anonymous):

yes,it will be -5 times sqrt 3

OpenStudy (anonymous):

-5?

OpenStudy (anonymous):

yeah,